OpenStudy (anonymous):

Solve for the variables: (r/3) - 5 = 63

OpenStudy (shamil98):

So solve for it... what do you think the first step is?

OpenStudy (shamil98):

You're main goal to solve for the variable is to get r completely by itself.

OpenStudy (anonymous):

Some of my answers are 22, 78, 174, 204

OpenStudy (judygreeneyes):

Multiply by 3 on both sides! r - 15 = 189 Then add 15 to both sides r = 189 + 15 = 204

OpenStudy (judygreeneyes):

Oooh -- got one!!

OpenStudy (anonymous):

Thx judygreeneyes

OpenStudy (anonymous):

Can u help me on this one solve x-1-(4x+3) < -19

OpenStudy (anonymous):

Some of my answers are x< -5, x<7,x>-7,x<5

OpenStudy (anonymous):

So do u know judygreeneyes

OpenStudy (judygreeneyes):

Here you go: Try to go back later and try to follow the steps, OK? It will help you on tests! x-1-(4x + 3) < -19 Do the parentheses first, in this case apply the negative sign to each term inside: x - 1 -4x - 3 < -19 Now combine the numbers and the x terms (simplify): -3x - 4 < -19 Move the 4 to the right side by adding 4 to both sides: -3x < -15 Now divide both sides by -3. When you do this (multiplying by a negative) in an inequality, you have to remember to switch the direction of the inequality: x > 5 You can plug it back in to be sure it is true!! For example, if x>5, then use x = 6: -3(6) < -16 -18<-15 true!!

OpenStudy (judygreeneyes):

Does this match any of your possible answers? If not, please check the question again, and I will re-check also.